Summary

The observation was more smooth than yesterday, almost no time wasted tonight. Seeing is good during the first half night (0.7-0.9 for red channel), and become bad during the second half night (1.1-1.4 for red channel) 2 EGS fields, 4 Bootes fields and 3 Herc fields are taken.

Details

Observers: Xiaohui Fan, Linhua Jiang, Fuyan Bian

Supporting Astronomer: Olga Kuhn

Operator: Aaron Ceranski

All Times UT

Second night of the AZ queue.

0238 The sky looks good, this is a little cloudy at the horizon. Not a photometric night.

0414 After struggling with the LBC for one and a half hours, we are taking data now. The estimated seeing is 1.0". The field is EGS_2a. The first image: the Y band focus is bad, and the U band seeing is 1.3". The second Y band image looks still bad, and we re-focus and re-do the whole OB

* 0431* We are re-doing EGS_2a, estimated seeing is 1.0". The first image: the Y band is a little improved, the seeing is 1.0" and the seeing for the U band is 1.1". Focus is being adjusted by the operator The second image: the Y band improved a lot, the seeing is 0.8" and the seeing for the U band is 1.1" The third one: The seeing for the Y band is 0.7" and the seeing for U band is 1.1" The forth one: can not find guide star for the red channel. The Y band image is trailed. The seeing for the U band is 1.1". The fifth one: The seeing for the Y band is 0.8" and for the U band is 1.1"

0458 We are moving to EGS_2b, the estimated seeing is 0.82". The first image: The seeing for the Y band is 0.8 and for the U band is 1.2"; The second one: The Y band image is bad (sometimes the guide star is very diffused) and seeing for the U band is 1.2"; The blue channel has a temperature problem, so the blue channel stops.; The third one: The seeing for the Y band is 0.8"; We have to stop to restart the LBC.

0533 We moving to b01r, estimated seeing is 0.7". The first image: The seeing is 0.7" for the Y band and the seeing changes much from the central region to the edge, from 0.7" to 0.9" The U band seeing is 1.1"; The second one: The Y band seeing is 0.8" and the U band seeing is 1.0"; The third one: The Y band seeing is 0.7" and the U band seeing is 0.9"; The forth one: It is similar to the third one; The fifth one: Both Y and U band images are bad. (lbcr.20080530.055409.fits and lbcb.20080530.055412.fits) telescope jumped.

0621 We moving to b03r, estimated seeing is 0.6". The first image: The Y band seeing is 0.8" and the U band seeing is 1.2", the telescope jumped a little; The second image: the telescope jumped, both channels images are bad; The third one: look better, the Y band seeing is 0.7" but bad at the edge (0.9"), and the U band seeing is 1"; The forth one: The seeing becomes a little worse, the Y band seeing is 0.8" and the U band seeing is 1.3"; The fifth one: The blue one fails and the Y band seeing is 0.8"; The blue channel met problems, we restarted the LBC.

0658 Moving to b05r, estimated seeing is 0.7". The first image: The Y band seeing is 0.9" the bottom conner for the first chip is very bad, and the U band seeing is 1.4"; The second image: Y band seeing is 0.8" and the U band seeing is 1.1". The telescope jumped a little; The third image: Y band seeing is 0.9" and the U band seeing is 1.3". The telescope jumped a little (bigger than the last one); The forth image: Y band seeing is 0.8" and the U band seeing is 1.4". The fifth image: Y band seeing is 0.8" and the U band seeing is 1.2"

0727 Moving to b10r, estimated seeing is 0.6". The first image: The Y band seeing is 0.8" and the U band seeing is 1.2"; The blue one failed. The second image: The telescope jumped. We have to stop We will redo the b10r field.

0747 Redoing the b10r field, estimated seeing is 0.6". The first image: The Y band seeing is 0.8" and the U band seeing is 1.3"; The second one : The Y band seeing is 0.8" and the U band seeing is 1.3"; The third one: The Y band seeing is 0.7 and the U band seeing is 1.2"; The fourth one: The telescope jumped The fifth one: The Y band seeing is 0.8" and the U band seeing is 1.1".

0821 Moving to the Herc_3 field, estimated seeing is 0.7" The first image: the r band seeing is 0.9" and B band seeing is 1.3" The second image: the r band seeing is 0.9" and B band seeing is 1.4", and the star shape in Band is bad; The third one: The r band seeing is 0.9" and B band seeing is 1.2", the star shape is a little improved; The forth one: The r band seeing is 0.9" and B band seeing is 1.1". The fifth one: The r band seeing is 0.9" and B band seeing is 1.4" The sixth one: The r band seeing is 1.0" and B band is 1.3" All the B band images is elongated. So we redo the B band image

0908 Redo the B band for the Herc_3 filed, estimated seeing is 1.2" for B band. The B band seeing is very stable, and it is around 1.3"

0942 Moving to the Herc_4 field, estimated seeing is 0.8 for r band. The first image: The r band seeing is 1.3" and B band seeing is 1.5". The star shape looks good; The second one: The r band seeing is 1.1" and B band seeing is 1.4". The star shape in B band image looks not very well; The third one: Seeing becomes bad. The r band seeing is 1.5" and B band seeing is 1.7"; The fourth one: The r band seeing is 1.2" and B band seeing is 1.3"; The fifth one: The r band seeing is 1.4" and B band seeing is 1.5" and the telescope jump a little (very tiny); The sixth one: The r band seeing is 1.2" and B band seeing is 1.7" and the band image is elongated.

10:24 Moving to the Herc_5 field, estimate seeing is 1.1 for r band. The first image: The r band seeing is 1.3" and the B band seeing is 1.7". The B band image is a little elongated. The second image is similar to the first one. The third one: The r band seeing is 1.2" and and B band seeing is 1.5". The forth one: The r band seeing is 1.2", and B band image is a little elongated and the seeing is 1.7"; The fifth one: The r band seeing is 1.2" and B band seeing is 1.3"; The sixth one: The telescope jumped.

11:00 Starting to take bias

11:25 Starting to take sky flat.

-- XiaohuiFan - 30 May 2008
Topic revision: r4 - 30 May 2008, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback